# Active pattern formation emergent from single-species nonreciprocity

### v1

@@ Abstract
Fig. 3: Dynamical processes of active pattern formation. Snapshots of density profile at different time stages of system evolution for two sample average densities $\bar{\rho}=0.7$ (a, c) and $\bar{\rho}=1$ (b, d), illustrating the rich dynamics of phase-separated fibers/strands which leads to the persistent variation of active branches (a, b with $\vartheta=110^{\circ}$, $90^{\circ}$; Supplementary Videos 6 and 7) or the self-knitting of active yarn (c, d with $\vartheta=130^{\circ}$; Supplementary Video 8).
Active branches and active yarn
One of our key findings is the emergence of self-migrating chiral patterns induced by single-species nonreciprocity, in the form of active branches and active yarn, at high enough opening angles or average particle densities. These patterns break time-reversal symmetry (traveling unidirectionally) and 2D parity symmetry (lacking mirror symmetry axes in space; Fig. 2b,c,d) for both particle and polarization density fields. The formation and nonlinear evolution of these exotic patterns are well beyond the stage of bifurcation and instability development, as illustrated in Fig. 3.
Figure 3 shows that two categories of pattern-forming dynamics can be identified. For intermediate vision-cone opening angles, we observe persistently varying active branches, while for large opening angles, the banding of active yarn occurs. For the dynamics of active branches there are two sub-types: At low enough average density (e.g., $\bar{\rho}=0.7$ for $\vartheta=110^{\circ}$; Supplementary Video 6), a microphase-separated stripe developed via instability sharpens to an individual fiber or thread, which splits and self-interweaves to form twisted or 2D double-helix type configurations. The strands then deform and bifurcate/branch, before breaking apart and constantly dispersing or regrouping. This double-twisting behavior does not occur at larger $\bar{\rho}$ (e.g., $\bar{\rho}=1$ for $\vartheta=90^{\circ}$; Supplementary Video 7), where one instead finds the curving and undulating of individual sharpened fibers during the secondary stage of phase separation. This is followed by their local self-braiding and bundling to form loose, varying clusters in the subsequent tertiary stage of evolution. For both sub-types the resulting branched traveling texture is highly dynamical, aggregating and dispersing incessantly.
For the second category – formation of active yarn at large opening angles – the branching occurs earlier and directly from the bent sharpening fibers. The multi-branched threads aggregate and merge to compact bands with fluctuating interfaces that separate from the homogeneous melt, yielding a self-knitting process of a strongly condensed, active yarn pattern (see Fig. 3 at $\vartheta=130^{\circ}$ and Supplementary Video 8). The pattern not only shows persistent variations of micro-textures within each band, but could also evolve via an alternating process of aggregation/breaking/rebanding with the change of traveling band orientation, while the overall pattern morphology still retains without coarsening with time (see Extended Data Fig. 1 and Supplementary Video 9).
Both active branches and active yarn patterns evolve at two different length and two separate time scales. The fine branched textures evolve rapidly and perpetually on short timescales, while the flow and variation of the entire pattern structure occur on longer timescales (see, e.g., Supplementary Videos 2-4, 9). As seen also in Extended Data Fig. 2 for the estimated probability distribution of velocity, the average migration velocity of the pattern, with value close to the sharp peak of the skewed distribution observed, corresponds to the overall timescale of pattern flow, while the extended range of the distribution, particularly for large magnitudes of velocity, reflects the fast variations of fine textures. Spatially, the local clustering during the variation of active branches, and more prominently the dense banding of active yarn, represent a large-scale phase separation between a surrounding homogeneous state and bundles or bands with fine-scale microphase-separated branched structures inside. Our simulations show that the larger length scale of bulk separation depends on the size of the simulation domain, suggesting that this is related to phase coexistence. The small length scale, on the other hand, depends on the physical parameters of the system and is about an order of magnitude smaller than the characteristic width of active yarn which is estimated in Extended Data Fig. 3. It is important to note that the formation of active yarn is beyond the traditional phase coexistence and is characteristic of a new dynamical pattern, since it involves the simultaneous development of small-scale multi-branching and larger-scale local bundling or banding (Fig. 3 and Supplementary Video 8), belonging to a different category of formation process as compared to active branches (noting also the difference of fine textures between active yarn with multi-branched strands within the highly condensed band and active branches consisting of mostly bifurcated fibers). Thus our results demonstrate the realizability of this unique type of active yarn pattern which self-organizes and evolves at multiple-scales through simultaneous microphase and bulk phase separations.
Microscopically, the banding phenomenon, particularly during the active yarn formation, can be understood by noting that, since the interaction combines a strong and short-ranged repulsive force with a weaker long-ranged attractive force (equation (9)), the interaction of a particle with its nearest neighbors is dominated by repulsion, whereas the interaction with its next-nearest neighbors is dominated by attraction. At very large vision-cone angle (e.g., $2\vartheta\geq 240^{\circ}$) with narrow blind spot, the net attraction from a broad range of next neighbors would be strong enough to forge and confine the particle, then bundling or banding together with other nearby particles while traveling concurrently. Thus narrower, denser bands can be expected for larger $\vartheta$ and smaller $\bar{\rho}$, consistent with the numerical simulation results shown in Extended Data Fig. 3. If the perception is further increased to $2\vartheta=2\pi$, there would be no blind spot and the interactions are reciprocal; the strong net repulsion would then push away particles from all directions ($\langle\mathbf{F}_{2\rightarrow 1}\rangle_{\rm local}=\mathbf{0}$ from equation (13)), making the density distribution disordered without any phase separation.
Active droplet growth dynamics
To further examine the chiral pattern dynamics, we simulate the growth process of an active droplet developing branched structures when embedded in a homogeneous flocking medium. As shown in Fig. 4, the self-traveling droplet domain grows slowly during the early-stage instability development of microphase separation to form weak stripes within, while the growth accelerates at intermediate time stage when the combination of microphase and bulk phase separations occurs during the secondary and tertiary processes of evolution; i.e., the stripes sharpen to become fibers which then curve/twist and branch, together with their local aggregation and bundling as well as segregating at a larger scale to the two internal ends of droplet, overall manifesting as a growing and moving “ball of wool”. A much faster growth takes place at later stage due to simulation boundary effects. These results are generic and consistent for different $\vartheta$ and $\bar{\rho}$ yielding active branched and yarn states.
Fig. 4: Active droplet growth. The droplet size $L(t)$ as a function of time $t$, at $\bar{\rho}=1$ and $\vartheta$ ranging from $90^{\circ}$ to $160^{\circ}$ for active branches or active yarn patterns (inset: $\bar{\rho}=1,1.2,1.5$ at $\vartheta=140^{\circ}$). The time-evolving snapshots, which are small portions of the large system simulated, correspond to the stars indicated on the $\vartheta=140^{\circ}$ curve. Also shown are the fittings to a power law scaling $L\propto t^{n}$ at early time (dashed lines), with $n=0.475\pm 0.001$ at $\vartheta=90^{\circ}$ and $n=0.435\pm 0.002$ at $\vartheta=140^{\circ}$, before the crossover to a linear regime of $L\propto t$.
It has been known [36] that the growth of droplet or domain size $L(t)$ obeys a power law behavior, $L\sim t^{n}$, with the scaling exponent $n=1/3$ for the diffusion-controlled mechanism (Lifshitz-Slyozov-Wagner law), $n=1/2$ for the curvature-driven growth, and $n=1$, i.e., linear growth, when the hydrodynamic advection effect dominates, while activity could possibly arrest phase separation and growth in “wet” systems [27, 37]. Our simulation results indicate a crossover between two scaling regimes before the influence of boundaries (Fig. 4), from an early-stage one showing only microphase separation with an exponent $n$ between $1/3$ and $1/2$, which can be attributed to a combination of diffusion-driven and curvature-driven effects, to a linear growth regime with $L\sim t$ when the growth mechanism is dominated by the combined phase separations induced by nonreciprocity. In contrast to the linear growth previously found in “wet” systems [27, 37], both advective effect and activity arrest are absent in the “dry” system studied here.
Conclusions
We have constructed a nonreciprocal continuum field theory, Active Model N, which incorporates the many-body effects induced by nonreciprocity and constitutes a minimal model for collective dynamics in the general class of single-species systems characterized by nonreciprocal forces. It allows us to reveal the intriguing results arising from the mechanism of competition between repulsive and attractive interactions of different strength and range and their combination with interparticle nonreciprocity and local alignment. This leads to the emergent phenomenon of single-species active pattern formation with PT symmetry breaking, a hallmark of nonreciprocal phase transitions. The resulting out-of-equilibrium chiral branched patterns, particularly the self-knitted active yarn, originate from a new class of active phase separation which simultaneously incorporates nonreciprocity-induced micro- and bulk phase separations, with local structural details being highly dynamical and varying persistently with time.
A similar behavior of banding and self-traveling of single-species agents with vision-cone interactions can be found in various real biological systems, such as the migration of caterpillars or ants [38] and the marching of penguins [39] or pedestrians [40]. The field theory developed here and our results of active pattern formation and dynamical mechanisms are expected to be generic and extendable to a range of nonreciprocal systems, particularly Janus colloids [3], that can be realized and controlled in experiments. Moreover, the active yarn phase, where a rather chaotic but patterned state is embedded within a homogeneous medium, has interesting potential applications in the realization of programmable materials. For example, if the particles are immersed in a fluid, this effect allows to achieve controlled mixing of this fluid in one region (the size of which can be controlled by tuning $\vartheta$ or $\bar{\rho}$) while leaving it unmixed elsewhere. If one additionally incorporates a mechanism in which the particles can adapt the value of $\vartheta$ in response to external stimuli, it would also allow for the realization of adaptive or even intelligent matter [41, 42].
Acknowledgements
We thank Filippo De Luca, Ken Elder, Helder Hugo, Jonas Mayer-Martins, and Simiso Mkhonta for helpful discussions. Z.-F.H. acknowledges support from the National Science Foundation under Grant No. DMR-2006446. M.t.V., R.W., and H.L. are funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-IDs 525063330, 433682494 – SFB 1459, and SPP 2265 – LO418/25.
Author contributions
Z.-F.H. and H.L. initiated and designed the research. M.t.V. conducted the theoretical derivation of continuum field equations. H.L., M.t.V., and R.W. developed the microscopic model. Z.-F.H. conducted the bifurcation analysis and performed numerical simulations. Z.-F.H., M.t.V., and H.L. wrote the manuscript. Z.-F.H., R.W., and H.L. prepared the figures. All authors analyzed the results, discussed the data, and revised the manuscript.
Competing interests
The authors declare no competing interests.
Extended Data Fig. 1: Time variation of active yarn pattern. During the nonreciprocity-induced self migration, the orientation and morphology of active yarn can vary with time, as illustrated in this sample simulation at $\bar{\rho}=1.5$ and $\vartheta=150^{\circ}$ (see also Supplemental Video 9).
Extended Data Fig. 2: Probability distribution of polarization and velocity. The probability density function (PDF) for the polarization component $P_{x}$ or $P_{y}$ is calculated from simulations at $\bar{\rho}=1.2$ and opening angle $\vartheta=90^{\circ}$ (active branches at $t=50000$) or $140^{\circ}$ (active yarn at $t=37500$, with $P_{x}=P_{\perp}$ and $P_{y}=P_{\parallel}$ in this example). It corresponds to the estimated probability distribution of locally coarse-grained particle velocity. The average migration velocity $\bar{\mathbf{v}}=(\bar{v}_{x},\bar{v}_{y})$, which is proportional to the reverse average polarization $-\langle\mathbf{P}\rangle$, is indicated in the inset images.
Extended Data Fig. 3: Width of active yarn as a function of opening angle $\vartheta$. Results are estimated from simulations at $\bar{\rho}=1.2$ and $1.5$ with system size $512\times 512$, showing the decrease of active yarn banding width with the increase of $\vartheta$ and the decrease of $\bar{\rho}$.

### v2

@@ C.2.2 Secondary instability: The nonreciprocal cases $0<\vartheta<\pi$
Appendix D Model parameters and numerical simulations
In this study, we use the soft-interaction form of Eq. 14 for the force and torque functions and choose the model parameters
$\beta aF_{0}=0.1,\;\frac{F_{1}}{F_{0}}=0.001,\;\beta b=0.03,\;\alpha=2,\;\frac{c}{a}=2.$ (141)
From Eqs. (90)–(94), the coupling coefficients $B_{i}$ in Active Model N are then found to be
$B_{1}$ $=\pi\beta aF_{0}\left(1-2\alpha^{2}\frac{F_{1}}{F_{0}}\right)\sin\vartheta=0.0992\pi\sin\vartheta,$
$B_{2}$ $=\frac{\pi}{2}\beta aF_{0}\left(\sqrt{\pi}-8\alpha^{3}\frac{F_{1}}{F_{0}}\right)\vartheta=0.05\pi(\sqrt{\pi}-0.064)\vartheta,$
$B_{3}$ $=\frac{\pi}{8}\beta aF_{0}\left(1-12\alpha^{4}\frac{F_{1}}{F_{0}}\right)\sin\vartheta=0.0101\pi\sin\vartheta,$
$B_{4}$ $=\frac{3}{32}\pi\beta aF_{0}\left(\sqrt{\pi}-64\alpha^{5}\frac{F_{1}}{F_{0}}\right)\vartheta$
$=-(3\pi/320)(2.048-\sqrt{\pi})\vartheta,$
$B_{5}$ $=-2\pi\beta b\left(\frac{c}{a}\right)^{2}\vartheta=-0.24\pi\vartheta,$ (142)
where we have rescaled $B_{1}\rightarrow B_{1}/a$, $B_{2}\rightarrow B_{2}/a^{2}$, $B_{3}\rightarrow B_{3}/a^{3}$, $B_{4}\rightarrow B_{4}/a^{4}$, and $B_{5}\rightarrow B_{5}/a^{2}$ in Eqs. (90)–(94) to make all the parameters dimensionless. In addition, $\tilde{D}_{\mathrm{R}}=a^{2}D_{\mathrm{R}}/D_{\mathrm{T}}=0.1$ is used.
Equations (11), (12), and (13) of Active Model N are solved numerically using a pseudospectral method with periodic boundary conditions. The simulated system size ranges from $256\times 256$ to $2048\times 2048$ grid points, with grid spacing $\Delta x=\Delta y=1$. The time step size $\Delta t$ is fixed within each simulation, but chosen from $0.001$ to $0.1$ depending on the values of $\bar{\rho}$ and $\vartheta$ to ensure numerical convergence and accuracy. The simulation snapshots shown in Figs. 2(a)–2(e) and Fig. 4 and results of Figs. 57 are for a system size $512\times 512$. For most simulation results presented here (all except for Fig. 2(e)), the initial conditions are chosen as a homogeneous or uniform flocking state with homogeneous particle density $\rho=\bar{\rho}$ and polarization $\mathbf{P}=\mathbf{\bar{P}}$ imposed by random initial fluctuations. Similar and consistent results are obtained if using different values of initial polarization alignment $\mathbf{\bar{P}}=(\bar{P}_{x},\bar{P}_{y})$. A different initial setup is used for the reciprocal limit of $\vartheta=\pi$ (Fig. 2(e)), where random initial conditions are adopted for both particle density and polarization fields in the fully disordered state with $\rho=\bar{\rho}$ and $\mathbf{P}=\mathbf{0}$.
All the results of active droplet growth dynamics given in Fig. 8 are obtained from simulations with a system size $2048\times 2048$. Initially both the droplet nucleus and outside medium are in the homogeneous flocking phase with $\rho=\bar{\rho}$ and $\mathbf{P}=\bar{\mathbf{P}}=(\bar{P},0)$ with $\bar{P}=|\bar{\mathbf{P}}|$ determined by Eq. 16, such that the droplet self-travels along the negative $x$ direction. Initial noise fluctuations are set inside the small circular nucleus of radius $r_{0}=5$, but not in the outside medium, which allows the droplet to expand and grow with time, with the development of patterns inside.